OpenStudy (4meisu):

Consider the expansion (x + 2)^11 What are the number of terms in this expansion. Find the term containing x^2

OpenStudy (anonymous):

Are you familiar with the binomial expansion?

OpenStudy (4meisu):

Our teacher briefly explained it and that was that..

OpenStudy (anonymous):

The binomial theorem says that: \[(x+y)^n = \sum_{k=0}^{n}\left(\begin{matrix}n \\ k\end{matrix}\right) x^{n-k} y^k\] So we have n=11, I'll give an example of how to find another term and you can use that to find the term with x^2. Say we want the term with x^4. Then we want the x^(n-k) to be x^4, so n-k=11-k=4, so k must be 7. So we find the term with n=11 and k=7. First we have the binomial co-efficient \[\left(\begin{matrix}11 \\ 7\end{matrix}\right)=330 \] (use the nCr button on the calculator) Then we will have x^(n-k)=x^(11-7)=x^4, and 2^k=2^7. So when we multiply all these parts together we get 330x^4 times 2^7 = 42,240x^4 as the term containing x^4. Are you able to apply this to find the term containing x^2? Please say if you don't understand something I've said or if I can go through the question step by step.

OpenStudy (4meisu):

Mm okay. How would you find the numbers in the expansion though?

OpenStudy (anonymous):

You have the three components for each term: -The x component, which will have power n-k. -The y component (or in this question, y=2 so we will have a power of 2), so for term k we will have 2^k. -The binomial coefficient. Since we know the value of y the numbers in the expansion will come from the binomial coefficient multiplied by the respective power of 2. Let me give another example: Say we want the term with x^7. Then n-k=7, so 11-k=7 and k=4. So for the numbers which multiply x^7, we will have: 2^k=2^4=16. The binomial coefficient 11C4=330. Multiplying these together we get the term 5,280x^7. Does this make more sense? Can you see where the numbers are coming from?

OpenStudy (4meisu):

Yeah, thank you :)

OpenStudy (anonymous):

Great :) What did you get as your x^2 term?